OpenStudy (anonymous):

find the value of the csc 40 degrees using your calculator

OpenStudy (anonymous):

yes, use your calculator

OpenStudy (anonymous):

first find \[\sin(40)\] then find the reciprocal

OpenStudy (anonymous):

1/sin(40)

OpenStudy (anonymous):

whats CSC?

OpenStudy (anonymous):

csc=1/sin sec=1/cos tan=sin/cos

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

156 in deg, use your calculator, i just used the calculator in my computer, try in your calcu